OpenStudy (anonymous):

f(x)=cx for x=1,2,...,n find c so that f(x) is a valid probability function , then derive the mean of X

OpenStudy (anonymous):

For this to be a valid PDF, you must satisfy \[\large\sum_{x=1}^n f(x)=\sum_{x=1}^ncx=1\]

OpenStudy (anonymous):

Recall the formula, \[\large\sum_{k=1}^nk=\frac{n(n+1)}{2}\] Using this gives \[\large\sum_{x=1}^ncx=c\sum_{x=1}^nx=\frac{cn(n+1)}{2}=1~~\iff~~c=\frac{2}{n(n+1)}\]

OpenStudy (anonymous):

Cheers for that SithsAndGiggles, have been stuck on that a while- need to do some catchup work ! appreciate it

OpenStudy (anonymous):

You're welcome! The part with the mean is fairly simple. The mean is given by the expected value formula, \[E(X)={\large\sum_{x=1}^nxf(x)}\] You might find this formula useful: \[\large\sum_{k=1}^nk^2=\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

sorry for my ignorance, but could you possibly show me how you would derive the mean? (will let me know if i did it correctly) thanks

OpenStudy (anonymous):

As per the formula for expected value, \[\large\begin{align*}E(X)&=\sum_{x=1}^nxf(x)\\ &=\sum_{x=1}^nx(cx)\\ &=c\sum_{x=1}^nx^2\\ &=c\frac{n(n+1)(2n+1)}{6}\\ &=\frac{2}{n(n+1)}\cdot\frac{n(n+1)(2n+1)}{6}\\ &=\frac{2n+1}{3}\\ \end{align*}\]